OpenStudy (anonymous):

Which of the following is the equation of a line that passes through the point (1,4) and is parallel to the x-axis? A. y = 4 B. x = 1 C. y = 1 D. x = 4

OpenStudy (shamil98):

Parallel lines have equal slopes. We assume the slope is zero for now. To find the y - intercept. Plug in the y value . 4 = 0(1) + b 4= b so the equation is y = 0x + 4 or y = 4

OpenStudy (shamil98):

The slope of the x-axis is zero forgot to mention that.

OpenStudy (anonymous):

oh okay that makes sense! so the answer is y=4?

OpenStudy (shamil98):

Yes.

OpenStudy (anonymous):

so every time its an ordered pair asking that kind of question the slope is automatically 0 unless it says otherwise

OpenStudy (shamil98):

No, it says parallel to the x-axis, parallel line have equal slopes, the slope of the x axis is 0, i made a mistake earlier in saying that.

OpenStudy (anonymous):

oh okay :)